# Mixed-state geometric phases of coherent and squeezed spin states

### v1

@@ III.1 The Uhlmann phase
Using Eq. (III.1), the Uhlmann phase generated in a cyclic process can be evaluated. The $A_{U}$ of the CSS only differs by a minus sign from that of the spin-$j$ system discussed in Ref. [66]. Therefore, their Uhlmann phases only differ by a minus sign. As explained in Sec. II, this is because the CSS is related to the spin-$j$ system with a particular Hamiltonian. Since the cases with $j=\frac{1}{2}$ and $1$ of the corresponding spin-$j$ systems have already been studied in Ref. [66], we will focus here on the $j=\frac{3}{2}$ CSS. In Appendix A, we show that (I) the Uhlmann phase reduces to the Berry phase associated with the ground state as $T\rightarrow 0$, thereby confirming the correspondence explained in Ref. [21], and (II) the $T\rightarrow 0$ expressions of the CSSs agree with those of Ref. [56] for pure states.
@@
We plot $\theta_{U}$ as a function of $T$ in Figure 1. In the low temperature limit, $\theta_{U}=\pi$, which indicates that the topological properties of the system change after a cyclic evolution along the equator. More specifically, the final purification (in terms of the fiber in the bundle language) is “antiparallel” to the initial purification. As temperature increases and crosses the first critical temperature $T^{1}_{c}\approx 0.321\omega_{0}$, the value of $\theta_{U}$ suddenly drops to 0, implying that a topological change occurs in the evolution along the equator on the parameter space. The final purification now becomes “parallel” to the initial purification. As temperature continues to rise, subsequent topological transitions occur at $T^{2}_{c}\approx 0.376\omega_{0}$ and $T^{3}_{c}\approx 0.493\omega_{0}$. In the high temperature limit, $\theta_{U}$ remains zero, meaning the associated Uhlmann holonomy is topologically trivial. This is reasonable since $\lim_{T\rightarrow 0}\rho=\frac{1}{4}1_{4\times 4}$. Therefore, the evolution loop $\rho(t)$ essentially becomes a single point with its horizontal lift $W(t)$ also reduced to a single point [66]. Hence, there is no difference between the initial and final fibers anymore. We mention that bosonic and fermionic coherent states at finite temperatures, in contrast, have been shown to have smooth behavior of the Uhlmann phase without any jumps [21].
@@ III.2 The interferometric geometric phase
A possible choice is $\tan^{2}\frac{\theta}{2}\mathrm{d}\phi=0$, which implies either $\theta=0$ or $\mathrm{d}\phi=0$. The former leads to $\zeta=0$, indicating that $D(\zeta)=1$, which is a trivial evolution. Therefore, we choose the latter and set $\phi$ to be constant. This represents a longitude on the unit-sphere parameter space shown by the big circle in Figure 2.
@@
In the bottom panel of Figure 3, we choose a special case with $\theta_{f}=\frac{3}{4}\pi$ to show the discontinuity of the IGP at $T_{c}=0.408\omega_{0}$. Similar to the case of the Uhlmann phase, this can be recognized as a geometric phase transition induced by temperature. The only difference is that, in general, the IGP is not directly related to the topological properties of the system. Furthermore, finite-temperature geometric phase transitions of the IGP seem to be less common than those of the Uhlmann phase. For a two-level Hermitian quantum system, a finite change in temperature in general does not induce any jump of the IGP [67, 37]. For a three-level system, a careful design of the evolution process can make it happen [37]. Our results of the CSS thus illustrate additional transitions of the IGP and show that the CSS is suitable for studying interesting properties of mixed-state geometric phases.
@@
IV GEOMETRICAL PHASES OF SSS
IV.1 Uhlmann Phase
For the SSS shown in Eq. (5), the parameter space is one-dimensional. Therefore, we need to determine the value of $\Theta$ to define a cyclic process for density matrices, which is essential for generating the Uhlmann phase. The density matrix of a system in thermal equilibrium governed by $H(\Theta)$ show in Eq. (9) is given by $\rho(\Theta)=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta)}$ with $Z=\operatorname{Tr}(\mathrm{e}^{-\beta H(\Theta)})$. By examining Eq. (6), we observe that the quadratic term of $J_{x}$ appears in the exponent, making it difficult to derive a general expression for the SSS. Therefore, for a concise illustration of the geometric phases of the SSS, we will focus our studies on a fixed value of $j$. Since the behavior of the IGP for a two-level system ($j=\frac{1}{2}$) is relatively simple [67], we will begin with $j=1$ for our analysis.
@@ IV.1 Uhlmann Phase
Figure 5: Uhlmann phase $\theta_{U}$ as a function of temperature $T$ on semi-log scale. The smooth curve indicates the absence of a transition in this case.
Nevertheless, numerical results can be obtained by using the Trotter-Suzuki approximation [68, 69] to tackle the path-ordered expression. Our findings are presented in Figure 5, where $\theta_{U}$ is plotted as a function of $T$ on semi-log scale. Unlike the Uhlmann phase of the CSS, the Uhlmann phase of the SSS exhibits no discontinuities as the temperature varies, indicating the absence of a temperature-induced topological phase transition. The results show that at finite temperature, the initial and final purifications no longer remain parallel, but the “angle” between them changes continuously with temperature without any abrupt transition from “parallel” to “antiparallel” as seen previously in the CSS. In the infinite temperature limit, $\theta_{U}=2\pi\equiv 0\mod 2\pi$ again. This is reasonable since $\lim_{T\rightarrow+\infty}\rho$ is proportional to the identity matrix, and the associated horizontal lift shrinks to a single point, resulting in $\lim_{T\rightarrow+\infty}\theta_{U}=0\mod 2\pi$ [24].
IV.2 The interferometric geometric phase
@@
$=\sum_{m=-j}^{j}\left\langle jm\right|\frac{\mathrm{e}^{-\beta\omega_{0}J_{z}}}{Z}\frac{2J_{-}J_{+}+2J_{z}}{8}\left|jm\right\rangle$
$=\sum_{m=-j}^{j}\frac{\mathrm{e}^{-m\beta\omega_{0}}}{Z}\frac{j^{2}+j-m^{2}}{4}.$ (63)
However, this condition cannot be satisfied at finite temperatures in general, so the IGP in its original form is not guaranteed. Therefore, the SSS provides an explicit example where a compatible evolution for generating the IGP is missing. This is in contrast to the Uhlmann phase based on the underlying Uhlmann bundle, which makes it possible to find the parallel transport on the geometric structure through the construction of the Uhlmann connection.
@@
Nevertheless, one can still obtain the non-adiabatic IGP based on Eq. (19) accumulated during an arbitrary evolution beyond the parallel-transport condition (15). To fairly compare with the results of the Uhlmann phase of the SSS, in the following we choose $j=1$ for the non-adiabatic IGP as well. The total phase of the $j=1$ SSS is given by
@@
In the top panel of Figure 6, the non-adiabatic IGP $\theta_{\text{G}}^{\prime}$ is plotted as a function of temperature $T$ and the squeezing parameter $\Theta_{f}\equiv\Theta(t_{\text{final}})$. Similar to the Uhlmann phase of the $j=1$ SSS, there is no temperature-induced geometric phase transition here. We show the behavior of $\theta^{\prime}_{\text{G}}$ as a function of $T$ for $\Theta_{f}=5.4$ in the bottom panel, in which $\theta^{\prime}_{\text{G}}$ continuously decreases with $T$. At finite temperatures, we find that $\theta^{\prime}_{\text{G}}$ also continuously varies with $\Theta_{f}$ without any discontinuities.

### v2

@@ III.1 The Uhlmann phase
Using Eq. (III.1), the Uhlmann phase generated in a cyclic process can be evaluated. The $A_{U}$ of the CSS only differs by a minus sign from that of the spin-$j$ system discussed in Ref. Hou et al. (2021b). Therefore, their Uhlmann phases only differ by a minus sign. As explained in Sec. II, this is because the CSS is related to the spin-$j$ system with a particular Hamiltonian. Since the cases with $j=\frac{1}{2}$ and $1$ of the corresponding spin-$j$ systems have already been studied in Ref. Hou et al. (2021b), we will focus here on the $j=\frac{3}{2}$ CSS. In Appendix A, we show that (I) the Uhlmann phase reduces to the Berry phase associated with the ground state as $T\rightarrow 0$, thereby confirming the correspondence explained in Ref. Wang et al. (2023), and (II) the $T\rightarrow 0$ expressions of the CSSs agree with those of Ref. Chaturvedi et al. (1987) for pure states.
@@
We plot $\theta_{U}$ as a function of $T$ in Figure 1. In the low temperature limit, $\theta_{U}=\pi$, which indicates that the topological properties of the system change after a cyclic evolution along the equator. More specifically, the final purification (in terms of the fiber in the bundle language) is “antiparallel” to the initial purification. As temperature increases and crosses the first critical temperature $T^{1}_{c}\approx 0.321\omega_{0}$, the value of $\theta_{U}$ suddenly drops to 0, implying that a topological change occurs in the evolution along the equator on the parameter space. The final purification now becomes “parallel” to the initial purification. As temperature continues to rise, subsequent topological transitions occur at $T^{2}_{c}\approx 0.376\omega_{0}$ and $T^{3}_{c}\approx 0.493\omega_{0}$. In the high temperature limit, $\theta_{U}$ remains zero, meaning the associated Uhlmann holonomy is topologically trivial. This is reasonable since $\lim_{T\rightarrow 0}\rho=\frac{1}{4}1_{4\times 4}$. Therefore, the evolution loop $\rho(t)$ essentially becomes a single point with its horizontal lift $W(t)$ also reduced to a single point Hou et al. (2021b). Hence, there is no difference between the initial and final fibers anymore. We mention that bosonic and fermionic coherent states at finite temperatures, in contrast, have been shown to have smooth behavior of the Uhlmann phase without any jumps Wang et al. (2023).
@@ III.2 The interferometric geometric phase
A possible choice is $\tan^{2}\frac{\theta}{2}\mathrm{d}\phi=0$, which implies either $\theta=0$ or $\mathrm{d}\phi=0$. The former leads to $\zeta=0$, indicating that $D(\zeta)=1$, which is a trivial evolution. Therefore, we choose the latter and set $\phi$ to be constant. This represents a longitude on the unit-sphere parameter space shown by the big circle in Figure 2. We remark that the same result can be obtained if the strengthened parallel-transport condition (19) is used instead.
@@
In the bottom panel of Figure 3, we choose a special case with $\theta_{f}=\frac{3}{4}\pi$ to show the discontinuity of the IGP at $T_{c}=0.408\omega_{0}$. Similar to the case of the Uhlmann phase, this can be recognized as a geometric phase transition induced by temperature. The only difference is that, in general, the IGP is not directly related to the topological properties of the system. Furthermore, finite-temperature geometric phase transitions of the IGP seem to be less common than those of the Uhlmann phase. For a two-level Hermitian quantum system, a finite change in temperature in general does not induce any jump of the IGP Andersson et al. (2016); Hou et al. (2023a). For a three-level system, a careful design of the evolution process can make it happen Hou et al. (2023a). Our results of the CSS thus illustrate additional transitions of the IGP and show that the CSS is suitable for studying interesting properties of mixed-state geometric phases.
@@
IV GEOMETRIC PHASES OF SSS
IV.1 One-Axis Squeezing
IV.1.1 Uhlmann Phase
For the SSS shown in Eq. (5), the parameter space is one-dimensional. Therefore, we need to determine the value of $\Theta$ to define a cyclic process for density matrices, which is essential for generating the Uhlmann phase. The density matrix of a system in thermal equilibrium governed by $H(\Theta)$ show in Eq. (9) is given by $\rho(\Theta)=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta)}$ with $Z=\operatorname{Tr}(\mathrm{e}^{-\beta H(\Theta)})$. By examining Eq. (6), we observe that the quadratic term of $J_{x}$ appears in the exponent, making it difficult to derive a general expression for the SSS. Therefore, for a concise illustration of the geometric phases of the SSS, we will focus our studies on a fixed value of $j$. Since the behavior of the IGP for a two-level system ($j=\frac{1}{2}$) is relatively simple Andersson et al. (2016), we will begin with $j=1$ for our analysis.
@@ IV.1.1 Uhlmann Phase
Figure 5: Uhlmann phase $\theta_{U}$ as a function of temperature $T$ for the $j=1$ one-axis SSS. A $\pi$-valued jump is observed at $T_{c}\approx 0.68\omega_{0}$.
Nevertheless, numerical results can be obtained by using the Trotter-Suzuki approximation Trotter (1959); Suzuki (1976) to tackle the path-ordered expression. Our findings are presented in Figure 5, where $\theta_{U}$ is plotted as a function of $T$. The Uhlmann phase of SSS, similar to that of CSS, exhibits a $\pi$-valued jump at a finite temperature $T_{c}\approx 0.68\omega_{0}$, indicating the existence of a temperature-induced topological phase transition. This result reveals that at finite temperatures, the initial and final purifications no longer remain parallel, and the “angle” between them undergoes an abrupt transition from “parallel” to “antiparallel”. As shown in Fig. 5, this explains the jump in the Uhlmann phase from $\pi$ to 00 at $T_{c}$. In the infinite-temperature limit, $\theta_{U}=0$ again. This is reasonable since $\lim_{T\rightarrow+\infty}\rho$ is proportional to the identity matrix, and the associated horizontal lift shrinks to a single point, resulting in $\lim_{T\rightarrow+\infty}\theta_{U}=0\mod 2\pi$ Hou et al. (2021a).
IV.1.2 The interferometric geometric phase
@@
$=\sum_{m=-j}^{j}\left\langle jm\right|\frac{\mathrm{e}^{-\beta\omega_{0}J_{z}}}{Z}\frac{2J_{-}J_{+}+2J_{z}}{8}\left|jm\right\rangle\dot{\Theta}$
$=\sum_{m=-j}^{j}\frac{\mathrm{e}^{-m\beta\omega_{0}}}{Z}\frac{j^{2}+j-m^{2}}{4}\dot{\Theta}.$ (66)
However, this condition cannot be satisfied at finite temperatures in general, so the IGP in its original form is not guaranteed. The parallel-transport condition (18) and its strengthened form (19) are both satisfied if $\dot{\Theta}=0$. However, this implies that $\Theta$ is a constant, such that $S(\Theta(t))=S(\Theta(0))=1$ for any $t$, resulting in a trivial evolution. Therefore, the IGP of the one-axis SSS also becomes trivial due to the parallel-transport condition.
@@
Nevertheless, one can obtain the non-adiabatic IGP based on Eq. (22) accumulated during an arbitrary evolution beyond the parallel-transport condition (18) or the strengthened version (19). To fairly compare with the results of the Uhlmann phase of the SSS, in the following we choose $j=1$ for the non-adiabatic IGP as well. The total phase of the $j=1$ SSS is given by
@@
In the top panel of Figure 6, the non-adiabatic IGP $\theta_{\text{G}}^{\prime}$ is plotted as a function of temperature $T$ and the squeezing parameter $\Theta_{f}\equiv\Theta(t_{\text{final}})$. Different from the Uhlmann phase of the $j=1$ SSS, there is no temperature-induced geometric phase transition here. We show the behavior of $\theta^{\prime}_{\text{G}}$ as a function of $T$ for $\Theta_{f}=5.4$ in the bottom panel, in which $\theta^{\prime}_{\text{G}}$ continuously decreases with $T$. At finite temperatures, we find that $\theta^{\prime}_{\text{G}}$ also continuously varies with $\Theta_{f}$ without any discontinuities.
@@
IV.2 Two-Axis Squeezing
Although realizing a nontrivial parallel transport process for the IGP via one-axis squeezing for the SSS is difficult, here we will show that it is possible in the two-axis squeezing scenario. In the following, we will investigate two-axis squeezing for the Uhlmann phase and IGP.
IV.2.1 Uhlmann Phase
For the SSS given by Eq. (10) under two-axis squeezing, the parameter space becomes two-dimensional. The Uhlmann process can be generated by a cyclic path either on the complex plane or equivalently on the sphere parametrized by $(\theta,\phi)$ with $z=\mathrm{e}^{-\mathrm{i}\phi}\tan\left(\frac{\theta}{2}\right)$. The density matrix of a system in thermal equilibrium governed by the continuous transformation $K(z(t))$ of the two-axis SSS at temperature $T$ then defines a loop $C(t):=\rho(z(t))$ in the manifold of density matrices. Explicitly,
$\rho(z)=\frac{1}{Z}\mathrm{e}^{-\beta\hat{H}(z)}=\frac{1}{Z}K(z)\mathrm{e}^{-\beta\hat{H}}K^{\dagger}(z),$ (69)
where $\hat{H}=\mathrm{e}^{-\beta\omega_{0}\hat{J}_{z}}$.
We continue to focus on the case of the $j=1$ SSS for comparison. The corresponding angular momentum operators are given in Eq. (55). From Eq. (11), we obtain
$K(\theta,\phi)=\left(\begin{array}[]{ccc}K_{11}&0&K_{13}\\ 0&1&0\\ -K_{13}&0&K_{11}\\ \end{array}\right),$ (73)
where
$K_{11}$ $=\cos\left[2\tan\left(\frac{\theta}{2}\right)\right],$
$K_{13}$ $=\frac{\mathrm{i}\mathrm{e}^{\mathrm{i}\phi}}{2}\left[\mathrm{e}^{2\mathrm{i}\tan\left(\frac{\theta}{2}\right)}-\mathrm{e}^{-2\mathrm{i}\tan\left(\frac{\theta}{2}\right)}\right].$ (74)
A relatively simple path can be chosen along the equator, i.e., $\theta=\frac{\pi}{2}$, where $\phi$ varies with time. We assume the evolution starts from $\phi=0$, so the initial density matrix is given by $\rho(0)=\frac{1}{Z}\mathrm{e}^{J^{2}_{+}-J^{2}_{-}}\mathrm{e}^{-\beta\omega_{0}J_{z}}\mathrm{e}^{J^{2}_{-}-J^{2}_{+}}$ according to Eq. (69).
The Uhlmann connection for the two-axis squeezing of the $j=1$ SSS can be obtained in similar way following that of the one-axis squeezing case. Based on Eq. (15),
$A_{U}=-\sum_{n\neq m}\chi_{nm}K\left|jn\right\rangle\left\langle jn\right|K^{\dagger}\mathrm{d}K\left|jm\right\rangle\left\langle jm\right|K^{\dagger}.$ (75)
However, the main difference lies in the complexity of $K(z)=\mathrm{e}^{zJ^{2}_{+}-\bar{z}J^{2}_{-}}$, which can be difficult to disentangle. As a result, obtaining an analytical expression for $K^{\dagger}\mathrm{d}K$ is nontrivial. Fortunately, since we are only interested in the case with $j=1$, $K^{\dagger}\mathrm{d}K$ can be represented by a $3\times 3$ matrix in the $|jn\rangle\langle jm|$ basis:
$K^{\dagger}\mathrm{d}K=\mathrm{i}\left(\begin{array}[]{ccc}-\sin^{2}(2)&0&-\frac{\mathrm{e}^{\mathrm{i}\phi}\sin(4)}{2}\\ 0&0&0\\ -\frac{\mathrm{e}^{-\mathrm{i}\phi}\sin(4)}{2}&0&\sin^{2}(2)\\ \end{array}\right)\mathrm{d}\phi.$ (79)
To compute $A_{U}$, a crucial step is to determine the coefficients $\chi_{nm}$. We find that all nonzero contributions to $A_{U}$ are governed by
$\chi\equiv\chi_{n,n+2}=\frac{\left(\mathrm{e}^{\frac{\beta\omega_{0}}{2}}-\mathrm{e}^{-\frac{\beta\omega_{0}}{2}}\right)^{2}}{\mathrm{e}^{\beta\omega_{0}}+\mathrm{e}^{-\beta\omega_{0}}},$ (80)
which follows from eliminating the diagonal elements of $K^{\dagger}\mathrm{d}K$. Eventually, we obtain the Uhlmann connection $A_{U}$ corresponding to the specific evolution path $C(t)$:
$A_{U}=\mathrm{i}\left(\begin{array}[]{ccc}-\frac{\sin^{2}(4)}{2}&0&\frac{\sin(8)e^{\phi}}{4}\\ 0&0&0\\ \frac{\sin(8)e^{-\mathrm{i}\phi}}{4}&0&\frac{\sin^{2}(4)}{2}\\ \end{array}\right)\mathrm{d}\phi.$ (84)
Figure 7: Uhlmann phase $\theta_{U}$ of two-axis squeezing of the $j=1$ SSS as a function of temperature $T$ on semi-log scale. The smooth curve indicates the absence of a transition.
By applying the Trotter-Suzuki approximation once again, our result of the Uhlmann phase as a function of temperature $T$ on a semi-logarithmic scale is shown in Figure 7. The smooth curve indicates that at finite temperatures, the “angle” between the initial and final purifications no longer jump between 00 and $\pi$ but varies continuously with temperature. This behavior is qualitatively different from the cases discussed earlier. Our results demonstrate that different quantum processes can exhibit different topological characteristics. The Uhlmann holonomy shown by the Uhlmann phase can be different even though the CSS, one-axis SSS, and two-axis SSS all follow unitary processes. In the infinite-temperature limit, we again find $\theta_{U}=2\pi\equiv 0\mod 2\pi$ for the two-axis squeezing, which is consistent with the Uhlmann phases of the CSS and one-axis SSS.
Figure 8: (Top panel) The IGP $\theta_{\text{G}}$ of the $j=1$ two-axis SSS as a function of $T$ and $\theta_{f}$, where $\theta_{f}\in[0,\frac{3\pi}{4}]$. $\theta_{\text{G}}$ is continuous with respect to both $\theta_{f}$ and $T$ (apart from a jump at $T=0$). (Bottom panels) $\theta_{\text{G}}$ vs $T$ at $\theta_{f}=\pi/2$, where $\theta_{\text{G}}$ exhibits no jumps (left), and $\theta_{\text{G}}$ vs $\theta_{f}$ at $T=0$, where there is a $\pi$-jump at $\theta_{f}\approx 1.33\,\text{rad}$ (right).
IV.2.2 The IGP
For the two-axis SSS, we find that the strengthened parallel transport condition (19) of the IGP holds for a nontrivial process which starts from the pole at $\theta=0$ and proceeds along the meridian with $\phi=\frac{\pi}{2}$ under the transformation $K(\theta,\phi)$. Of course, such a path also respects the condition (18). The density matrix $\rho(t)$ of this unitary evolution is
$\rho(z(t))=K(z(t))\rho(0)K^{\dagger}(z(t)),$ (85)
where $\rho(0)=\frac{\mathrm{e}^{-\beta\omega_{0}J_{z}}}{Z}$. In the $j=1$ representation, the matrix form of $K^{\dagger}\mathrm{d}K$ is
$K^{\dagger}\mathrm{d}K=-\mathrm{i}\left(\begin{array}[]{ccc}0&0&\frac{2}{\cos(\theta)+1}\\ 0&0&0\\ \frac{2}{\cos(\theta)+1}&0&0\\ \end{array}\right)\mathrm{d}\theta,$ (89)
where all diagonal elements are zero. Therefore, the parallel transport condition  (19) is equivalent to
$\langle jm|K^{\dagger}(t)\dot{K}(t)|jm\rangle=0,\quad m=-1,0,1,$ (90)
and is always satisfied. The weaker parallel transport condition (18) is also satisfied. In this case, the IGP is given by
$\theta_{\text{G}}(t)$ $=\arg\operatorname{Tr}[\rho(0)K(t)]$
$=\arg\operatorname{Tr}[\frac{1}{Z}\mathrm{e}^{-\beta\omega_{0}J_{z}}\mathrm{e}^{zJ^{2}_{+}-\bar{z}J^{2}_{-}}]$
$=\arg\left[\frac{2\cos\left(2\tan\left(\frac{\theta}{2}\right)\right)}{\text{sech}(\beta\omega_{0})+2}+\frac{1-2\mathrm{i}\sin\left(2\tan\left(\frac{\theta}{2}\right)\right)}{2\cosh(\beta\omega_{0})+1}\right].$ (91)
Since $\tan\left(\frac{\theta}{2}\right)\to\infty$ as $\theta\to\pi$ causes rapid oscillations in the values of $\theta_{\text{G}}(t)$, the range of $\theta$ is restricted to $[0,\frac{3}{4}\pi]$ in our numerical calculation to ensure accurate results. In the top panel of Figure 8, the IGP $\theta_{\text{G}}$ is plotted as a function of temperature $T$ and the two-axis squeezing parameter $\theta_{f}\equiv\theta(t_{\text{final}})$. In this two-axis squeezing model, similar to the Uhlmann phase, there is no temperature-induced geometric phase transition. In the bottom panel (left side), we show the behavior of $\theta_{\text{G}}$ as a function of $T$ for $\theta_{f}=\pi/2$, where $\theta_{\text{G}}$ continuously increases with $T$. At finite temperatures, similar to the non-adiabatic IGP of the one-axis squeezing of the SSS, we find that $\theta_{\text{G}}$ also changes continuously with $\theta_{f}$ without any abrupt jumps.
However, in the zero-temperature limit, $\theta_{\text{G}}=\arg\left[\cos\left(2\tan\left(\frac{\theta}{2}\right)\right)\right]$, which suggests that a $\pi$-jump in the IGP may occur at the zeros of $\cos\left(2\tan\left(\frac{\theta}{2}\right)\right)$, corresponding to sign changes within the function’s argument. Therefore, as $\theta_{f}$ crosses $\theta_{f}=2\tan^{-1}\left(\frac{\pi}{4}\right)\approx 1.33\,\text{rad}$, the value of $\theta_{\text{G}}$ undergoes a $\pi$-jump (from $2\pi$ to $\pi$ or vice versa), signaling a geometric phase transition. This behavior of $\theta_{\text{G}}$ is shown in the bottom right panel of Figure 8. For the $j=1$ two-axis SSS, however, no finite-temperature jumps occur for the mixed-state geometric phases studied here.
Figure 9: Contour plot of the non-adiabatic IGP $\theta^{\prime}_{\text{G}}$ for the two-axis SSS with $j=1$ as a function of temperature $T$ and the squeezing parameter $\theta_{f}$, where $\theta_{f}=\phi_{f}$ corresponding to $\theta_{f}\in[0,\frac{3\pi}{4}]$.
Similar to the previous examples, the non-adiabatic IGP of two-axis squeezing can be expressed as
$\theta^{\prime}_{\text{G}}(t)=\arg\operatorname{Tr}[\rho(0)K(t)]+\mathrm{i}\int_{0}^{t}\operatorname{Tr}\left[\rho(0)K^{\dagger}(t^{\prime})\dot{K}(t^{\prime})\right]\mathrm{d}t^{\prime}.$ (92)
The evolution is characterized by a simple curve $\kappa(t):\theta(t)=\phi(t)$ with $\theta_{f}\in[0,\frac{3\pi}{4}]$ selected in the parameter space. This curve starts at $\theta=0$ when $t=0$ and ends at any point $(\theta_{f},\phi_{f})$ along the path, where $\theta_{f}\equiv\theta(t_{\text{final}})$ and $\phi_{f}\equiv\phi(t_{\text{final}})$. To contrast with the IGP, we present the contour plot of $\theta^{\prime}_{\text{G}}$ as a function of $T$ and $\theta_{f}$ for the $j=1$ two-axis SSS. The overall configuration of this plot demonstrates remarkable similarity to Fig. 8. This is expected given that the second term in Eq. (92) represents a continuous function $g(\theta_{f},T)$ dependent on $\theta_{f}$ and temperature $T$. As shown in Figure 9, the non-adiabatic IGP $\theta^{\prime}_{\text{G}}$ remains continuous as a function of $T$, independent of the specific value of $\theta_{f}$.
Nevertheless, there is a $\pi$-jump of the non-adiabatic IGP at $\theta_{f}=1.33$ rad when $T=0$, similar to the IGP shown in Fig. 8. The evolution of the IGP, however, is generated by the path with constant $\phi$ and respects the parallel-transport condition. The discontinuity stems solely from the first term of Eq. (92), a feature that remains invariant irrespective of the parallel-transport condition. Consequently, the zero-temperature jump survives in the non-adiabatic IGP as well. These findings support feasibility for experimentally probing geometric phase transitions of mixed states. Even when evolution strictly adhering to the parallel-transport condition is experimentally challenging, the framework of non-adiabatic IGP may still characterize possible geometric properties of the system.